Pablo, Faustina, and Mark swim at a public pool each day for exercise. One day Pablo swam 5 more than 3 times as many laps as Mark, and Faustina swam four times as many laps as Mark. If the men swam 29 laps altogether, how many laps did each one swim?
What I have done:
29=5+3m
24=3m
24/3=3m/3
7=m
Pablo swam 7 laps
29=4m
29/4=4m/4
7 1/4=m
Faustina swam 7 and 1/4 laps
29-7=22-7 1/4=14 3/4
Mark swam 14 and 3/4 laps

Answer by rapaljer(4671)

Let m= Number of laps for Mark
3m+5 = Number of laps for Pablo
4m= Number of laps for Faustina

The TOTAL number of laps (that means add them up!) = 29

m+3m+5 + 4m=29
8m+5 = 29
8m+5-5 = 29-5
8m= 24
m=3 Mark
3m+5 = 3*3 + 5 = 14 Pablo
4m= 4*3= 12 = Faustina

Check: 3+14+12 = 29
It checks!!

For additional explanation on WORD PROBLEMS, please see my own website. To find my website, click on my tutor name "rapaljer" anywhere in algebra.com. Click on the link there that will take you to the Homepage of my website, which is a subwebsite at Seminole State College of Florida. From my "Homepage", look for the link "Basic, Intermediate, and College Alegebra: One Step at a Time." Click on "Basic Algebra", and look in "Chapter 1" for Sections 1.09 and 1.10 on Word Problems. Here you will find my own complete non-traditional, user-friendly explanation of the topics that my own students found MUCH easier to understand than the traditional textbooks! Many of my exercises here are explained in great detail in the MATH IN LIVING COLOR pages that go with this curriculum. Best of all, it's all FREE!

I have explanations on nearly every topic in algebra and many from geometry as well, and many videos from my younger days as well. Everything is FREE, and I don't even have advertising on the pages!! If you don't need help with math, then visit my website to see the JOKE PAGE--all clean jokes, safe for the little ears in your family!!!

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida